# When and How: Learning Identifiable Latent States for Nonstationary Time Series Forecasting

### v1

@@ 4.2 Modular Architectures with Stationary and Nonstationary Priors
Since we assume that the noise term in Equation (2) is independent with ${\mathbf{z}}_{t-1}^{s}$, we enforce the independence of the estimated noise $\hat{\epsilon}_{t}^{s}$ and we further have:
$\small\log p(\hat{{\mathbf{z}}}^{s}_{t}|{\mathbf{z}}^{s}_{t-1})=\log p(\hat{\epsilon}^{s}_{t})+\sum_{i=n_{d}+1}^{n}\log|\frac{\partial r_{i}^{s}}{\partial\hat{z}^{s}_{t-1,i}}|.$ (12)
@@
$\small p(\hat{{\mathbf{z}}}_{1:t}^{s})=p(\hat{{\mathbf{z}}}^{s}_{1})\prod_{\tau=2}^{t}\!\!\left(\sum_{i=n_{d}+1}^{n}\log p(\hat{\epsilon}^{s}_{\tau,i})+\!\!\!\!\sum_{i=n_{d}+1}^{n}\!\!\!\log|\frac{\partial r_{i}^{s}}{\partial\hat{z}^{s}_{\tau-1,i}}|\right),$ (13)
where $p(\hat{\epsilon}^{s}_{i})$ follow Gaussian distributions. And another prior $p(\hat{{\mathbf{z}}}_{t+1:T}^{s}|\hat{{\mathbf{z}}}_{1:t}^{s})$ follows a similar derivation.
Modular Architecture for the Nonstationary Prior Estimation. We employ a similar derivation and let $\{r_{i}^{e}\}$ be a set of learned inverse transition functions, which take the estimated environment labels $\hat{{\mathbf{e}}}_{t}$ and $\hat{{\mathbf{z}}}^{e}_{t}$ as input and output the noise term, i.e. $\hat{\epsilon}^{e}_{t}=r_{i}^{e}(\hat{{\mathbf{e}}}_{t},\hat{z}^{e}_{t,i})$.
Leaving $r_{i}^{e}$ be an MLP, we further devise another transformation $\phi^{e}:=\{\hat{{\mathbf{e}}}_{t},\hat{{\mathbf{z}}}^{e}_{t}\}\rightarrow\{\hat{{\mathbf{e}}}_{t},\hat{\epsilon}^{e}_{t}\}$ with its Jacobian is $\small{\mathbf{J}_{\phi^{e}}=\begin{pmatrix}\mathbb{I}&0\\ *&\text{diag}\left(\frac{\partial r_{i}^{e}}{\partial\hat{z}_{t,i}^{e}}\right)\end{pmatrix}}$, where *** denotes a matrix. Similar to the derivation of stationary prior, we have:
@@
$\small\begin{split}\ln p({\mathbf{z}}^{e}_{1:t})&=\mathbb{E}_{q(\hat{{\mathbf{e}}}_{1:t})}\sum_{\tau=1}^{t}\left(\sum_{i=1}^{n_{e}}\ln p(\hat{\epsilon}^{e}_{\tau,i})+\sum_{i=1}^{n_{e}}\ln|\frac{\partial r_{i}^{e}}{\partial\hat{z}^{e}_{\tau,i}}|\right).\end{split}$ (15)
Note that $q(\hat{{\mathbf{e}}}_{1:t})$ denotes the environment estimation module, which is implemented with an autoregressive hidden Markov model and generates latent environment indices with the help of the Viterbi algorithm (Song et al., 2023; Elliott et al., 2012). To optimize the autoregressive hidden Markov model, we need to maximize its free energy lower bound, which is shown as follows:
@@ 4.3 Model Summary
Considering that the autoregressive hidden Markov model converges much faster than the sequential variational inference model, we employ a two-phase training strategy. Specifically, we first minimize $\mathcal{L}_{HMM}$ to train the autoregressive hidden Markov model. Sequentially, we minimize $\mathcal{L}_{EBLO}$ by fixing the parameters of the autoregressive hidden Markov model. Since we use the historical observations ${\mathbf{x}}_{1:t}$ to generate $\hat{{\mathbf{e}}}_{1:t}$, which can be used to estimate the transition matrix $\hat{\mathbf{A}}$, during testing phase, we can estimate $\hat{{\mathbf{e}}}_{t+1:T}$ by sampling from $\hat{\mathbf{A}}$ as shown in the environment prediction block in Figure 3.
4.4 Complexity Analysis
We further provide time complexity for the IDEA model. We first let $L$ be the length of the time series and the model complexity depends on the autoregress hidden Markov model and the sequential variational inference model. First, suppose that $\bm{\mathrm{E}}$ is the number of latent environments, then the complexity of the autoregress hidden Markov model is $\mathcal{O}(\bm{\mathrm{E}}L)$. Moreover, suppose that $n$ is the number of latent variables, then the complexity of the sequential variational inference model is $\mathcal{O}(nL)$. In practice, the number of latent environments and the dimension of latent variables are low, so our model enjoys an efficient computation. Please refer to Appendix G for model efficiency comparison from the perspectives of performance, efficiency, and GPU memory.
@@ 5.1.1 Experimental Setup
Data Generation. We generate the simulated nonstationary time series data with 3 environments as mentioned in Figure 2 and Equation (1)-(3). Specifically, we first randomly initialized a Markov Chain with a transition matrix A. Sequentially, for each environment, we consider different Gaussian distributions and generate the nonstationary latent variables ${\mathbf{z}}_{t}^{e}$. As for the stationary latent variables, we employ an MLP with a LeakyReLU unit as the transition function. Note that we generate Dataset A and Dataset B with different time lag dependencies. Finally, we use a randomly initialized MLP to generate the observation data.
Evaluation Metrics. We consider three different metrics to evaluate the effectiveness of our method. First, to evaluate the identifiability of the stationary and nonstationary latent variables, consider the Mean Correlation Coefficient (MCC) on the test dataset, which is a standard metric for nonlinear ICA. A higher MCC denotes a better identification performance the model can achieve. Second, to evaluate the identifiability of the transition matrix A, we further consider the Mean Square Error (MSE) between the ground truth A and the estimated one. A lower value of MSE implies the model can identify the transition matrix better. Finally, we also consider the accuracy of ${\mathbf{e}}_{t}$ estimation since it reflects the performance of our model in detecting when the temporal distribution shift occurs. Please refer to Appendix F.1 for a detailed discussion about evaluation metrics.
Baselines. Besides the standard BetaVAE (Higgins et al., 2016) that does not consider any temporal and environment information, we also take some conventional nonlinear ICA methods into account like TCA (Hyvarinen & Morioka, 2016)and i-VAE (Khemakhem et al., 2020a). Moreover, we consider TDRL (Yao et al., 2022), which considers stationary and nonstationary causal dynamics, but requires observed environment variables. Finally, we consider the HMNLICL (Hälvä & Hyvarinen, 2020) and NCTRL (Song et al., 2023), which assume that the environment variables are latent but follow a Markov Chain.
@@ 5.1.2 Results and Discussion
Table 1: Experiment results of two synthetic datasets on baselines and proposed IDEA.
@@
BetaVAE (Higgins et al., 2016) 64.2 63.2 63.7
TCL (Hyvarinen & Morioka, 2016) 56.0 65.8 60.9
i-VAE (Khemakhem et al., 2020a) 76.9 73.0 74.9
HMNLICA (Hälvä & Hyvarinen, 2020) 83.2 74.5 78.8
TDRL (Yao et al., 2022) 78.5 78.8 78.6
NCTRL (Song et al., 2023) 81.4 79.4 80.4
IDEA 97.5 92.7 95.1
@@
We generate two different synthetic datasets with different time lags dependencies. Experiment results of MCC are shown in Table 1, and the experiment results of environment estimation accuracy and MSE can be found in Appendix F.1. According to the experiment results, we can draw the following conclusions: 1) We can find that the accuracies of environment estimation are high in both two datasets. Since the Dataset B contains more complex temporal relationships, the corresponding accuracy is slightly lower. 2) We can also find that the proposed IDEA model can reconstruct the latent variables under unknown temporal distribution shift with ideal MCC performance, i.e., $(>0.95)$ on average. In the meanwhile, the other compared methods, that do not use historical dependency, can hardly perform well. Moreover, the TDRL, which considers the temporal causal relationship, cannot obtain an ideal MCC performance since it requires observed environments. 3) Although baselines like HMNLICA and NCTRL assume latent environments on a Markov Chain, they perform worse than our IDEA. This is because both HMNLICA and NCTRL do not consider the stationary and nonstationary latent variables in the data generation process simultaneously.
Models IDEA Koopa SAN DLinear N-Transformer RevIN MICN TimeNets WITRAN
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 36-12 0.291 0.345 0.323 0.370 0.624 0.522 0.330 0.368 0.387 0.409 0.372 0.408 0.292 0.358 0.350 0.389 0.329 0.375
72-24 0.300 0.353 0.305 0.356 0.319 0.376 0.312 0.355 0.469 0.449 0.402 0.425 0.324 0.378 0.343 0.390 0.340 0.388
144-48 0.338 0.38 0.353 0.388 0.367 0.406 0.351 0.380 0.548 0.507 0.458 0.445 0.396 0.418 0.376 0.407 0.385 0.419
216-72 0.367 0.388 0.379 0.404 0.462 0.457 0.372 0.394 0.606 0.536 0.494 0.468 0.396 0.429 0.409 0.430 0.410 0.436
ETTh2 36-12 0.141 0.236 0.142 0.240 0.160 0.271 0.161 0.269 0.156 0.256 0.177 0.274 0.148 0.255 0.152 0.251 0.143 0.248
72-24 0.173 0.260 0.185 0.271 0.343 0.413 0.187 0.285 0.216 0.301 0.266 0.340 0.183 0.280 0.194 0.284 0.180 0.276
144-48 0.233 0.306 0.243 0.313 0.266 0.337 0.240 0.321 0.309 0.378 0.380 0.417 0.253 0.335 0.270 0.336 0.256 0.329
216-72 0.262 0.327 0.283 0.344 0.301 0.367 0.287 0.356 0.396 0.426 0.442 0.457 0.294 0.362 0.307 0.363 0.312 0.368
Exchange 36-12 0.014 0.074 0.015 0.079 0.014 0.074 0.016 0.088 0.017 0.085 0.022 0.102 0.018 0.093 0.016 0.084 0.014 0.075
72-24 0.023 0.102 0.027 0.113 0.024 0.105 0.026 0.110 0.031 0.122 0.045 0.152 0.027 0.115 0.031 0.123 0.025 0.107
144-48 0.042 0.141 0.050 0.156 0.046 0.153 0.049 0.160 0.077 0.196 0.127 0.261 0.045 0.153 0.072 0.194 0.046 0.148
216-72 0.065 0.180 0.075 0.193 0.066 0.186 0.071 0.189 0.123 0.252 0.233 0.357 0.064 0.183 0.115 0.251 0.071 0.185
ILI 36-12 1.218 0.694 1.994 0.880 2.296 1.027 2.620 1.154 1.491 0.757 2.637 1.094 4.847 1.570 2.406 0.840 2.138 0.914
72-24 1.680 0.809 2.077 0.919 2.472 1.074 2.733 1.177 2.551 1.039 2.653 1.116 4.776 1.556 2.270 0.988 2.867 1.080
108-36 1.792 0.869 1.836 0.881 3.228 1.209 2.271 1.094 2.227 1.018 2.696 1.139 4.917 1.584 2.978 1.123 3.147 1.151
144-48 1.883 0.926 2.072 0.941 3.842 1.280 2.490 1.187 2.595 1.081 2.960 1.167 4.804 1.584 2.696 1.098 3.388 1.232
weather 36-12 0.072 0.09 0.076 0.098 0.079 0.121 0.080 0.127 0.077 0.100 0.095 0.127 0.076 0.120 0.081 0.106 0.093 0.144
72-24 0.098 0.130 0.100 0.133 0.109 0.160 0.113 0.172 0.110 0.145 0.134 0.193 0.099 0.150 0.110 0.151 0.143 0.203
144-48 0.115 0.158 0.125 0.164 0.128 0.181 0.130 0.193 0.138 0.187 0.175 0.237 0.127 0.186 0.131 0.178 0.202 0.262
216-72 0.136 0.187 0.139 0.187 0.148 0.210 0.147 0.215 0.174 0.225 0.225 0.287 0.149 0.213 0.149 0.202 0.271 0.319
ECL 36-12 0.114 0.216 0.135 0.245 0.133 0.246 0.166 0.254 0.134 0.242 0.136 0.254 0.250 0.338 0.128 0.236 0.147 0.272
72-24 0.121 0.220 0.129 0.236 0.139 0.252 0.166 0.252 0.14 0.246 0.144 0.257 0.258 0.342 0.134 0.242 0.155 0.278
144-48 0.122 0.224 0.136 0.242 0.148 0.259 0.151 0.245 0.155 0.260 0.163 0.275 0.271 0.353 0.149 0.256 0.172 0.291
216-72 0.131 0.232 0.142 0.246 0.156 0.264 0.144 0.242 0.169 0.274 0.175 0.287 0.279 0.357 0.166 0.271 0.183 0.300
Traffic 36-12 0.457 0.300 0.472 0.312 0.489 0.304 0.602 0.381 0.626 0.319 0.567 0.373 0.583 0.312 0.609 0.306 0.802 0.387
72-24 0.458 0.305 0.450 0.309 0.495 0.310 0.613 0.386 0.574 0.310 0.548 0.363 0.618 0.334 0.553 0.296 0.811 0.401
144-48 0.410 0.283 0.420 0.302 0.489 0.309 0.472 0.319 0.590 0.330 0.570 0.361 0.616 0.327 0.555 0.301 0.830 0.406
216-72 0.403 0.277 0.424 0.308 0.500 0.312 0.441 0.300 0.604 0.338 0.547 0.336 0.650 0.343 0.577 0.311 0.854 0.415
@@ 5.2.1 Experiment Setup
Datasets. To evaluate the effectiveness of our IDEA method in real-world scenarios, we conduct experiments on eight real-world benchmark datasets that are widely used in nonstationary time series forecasting: ETT (Zhou et al., 2021), Exchange (Lai et al., 2018), ILI(CDC), electricity consuming load (ECL), weather (Wetterstation), traffic and M4 (Makridakis et al., 2020). More detailed descriptions of the datasets can be referred to Appendix F.2.1. We employ the same data preprocessing and split ratio in TimeNet (Wu et al., 2022). Similarly to Koopa (Liu et al., 2023b), for each forecasting window length $H$, we let the length of the lookback window $3H$ to make sure that each type of environment exists in the lookback window. Moreover, for each dataset, we consider different forecast lengths $H\in\{12,24,48,72\}$.
@@
Baselines. To evaluate the effectiveness of the proposed IDEA, we consider the following state-of-the-art deep forecasting models for time series data. First, we consider the methods for long-term forecasting including the TCN-based methods like TimesNet (Wu et al., 2022) and MICN (Wang et al., 2022), the MLP-based methods like DLinear (Zeng et al., 2023), as well as the recently proposed WITRAN (Jia et al., 2023). Moreover, we further consider the methods with the assumption that the temporal distribution shift occurs among instances like RevIN (Kim et al., 2021) and Nonstationary Transformer (Liu et al., 2022). Finally, we compare the nonstationary forecasting methods with the assumption that temporal distribution shift occurs uniformly in each instance, like Koopa (Liu et al., 2023b) and SAN (Liu et al., 2023c). Since SAN is a framework that boosts other forecasting models, we report the results of SAN+autoformer (Wu et al., 2021). We repeat each experiment over 3 random seeds and publish the average performance.
Experiment Details. We use ADAM optimizer (Kingma & Ba, 2014) in all experiments and report the mean squared error (MSE) and the mean absolute error (MAE) as evaluation metrics. All experiments are implemented by Pytorch on a single NVIDIA RTX A5000 24GB GPU.

### v4

@@ 4.2 Stationary and Nonstationary Priors Estimation
Since we assume that the noise term in Equation (2) is independent of ${\mathbf{z}}_{t-1}^{s}$, we enforce the independence of the estimated noise $\hat{\epsilon}_{t}^{s}$ and we further have:
$\small\log p(\hat{{\mathbf{z}}}^{s}_{t}|\hat{{\mathbf{z}}}^{s}_{t-1})=\log p(\hat{\epsilon}^{s}_{t})+\sum_{i=n_{d}+1}^{n}\log|\frac{\partial r_{i}^{s}}{\partial\hat{z}^{s}_{t,i}}|.$ (12)
@@
$\small p(\hat{{\mathbf{z}}}_{1:t}^{s})=p(\hat{{\mathbf{z}}}^{s}_{1})+\prod_{\tau=2}^{t}\!\!\left(\sum_{i=n_{d}+1}^{n}\log p(\hat{\epsilon}^{s}_{\tau,i})+\!\!\!\!\sum_{i=n_{d}+1}^{n}\!\!\!\log|\frac{\partial r_{i}^{s}}{\partial\hat{z}^{s}_{\tau,i}}|\right),$ (13)
where $p(\hat{\epsilon}^{s}_{i})$ follow Gaussian distributions. Another prior $p(\hat{{\mathbf{z}}}_{t+1:T}^{s}|\hat{{\mathbf{z}}}_{1:t}^{s})$ follows a similar derivation.
Modular Architecture for the Nonstationary Prior Estimation. We employ a similar derivation and let $\{r_{i}^{e}\}$ be a set of learned inverse transition functions, which take the estimated environment labels $\hat{{\mathbf{e}}}_{t}$ and $\hat{{\mathbf{z}}}^{e}_{t}$ as input and output the noise term, i.e. $\hat{\epsilon}^{e}_{t}=r_{i}^{e}(\hat{{\mathbf{e}}}_{t},\hat{z}^{e}_{t,i})$. Leaving $r_{i}^{e}$ be an MLP, we further devise another transformation $\phi^{e}:=\{\hat{{\mathbf{e}}}_{t},\hat{{\mathbf{z}}}^{e}_{t}\}\rightarrow\{\hat{{\mathbf{e}}}_{t},\hat{\epsilon}^{e}_{t}\}$ with its Jacobian $\small{\mathbf{J}_{\phi^{e}}=\begin{pmatrix}\mathbb{I}&0\\ *&\text{diag}\left(\frac{\partial r_{i}^{e}}{\partial\hat{z}_{t,i}^{e}}\right)\end{pmatrix}}$, where * denotes a matrix. Similarly, we have:
@@
$\small\begin{split}\ln p(\hat{{\mathbf{z}}}^{e}_{1:t})&=\mathbb{E}_{q(\hat{{\mathbf{e}}}_{1:t})}\sum_{\tau=1}^{t}\left(\sum_{i=1}^{n_{e}}\ln p(\hat{\epsilon}^{e}_{\tau,i})+\sum_{i=1}^{n_{e}}\ln|\frac{\partial r_{i}^{e}}{\partial\hat{z}^{e}_{\tau,i}}|\right).\end{split}$ (15)
Note that $q(\hat{{\mathbf{e}}}_{1:t})$ denotes the environment estimation module, which is implemented with an autoregressive hidden Markov model and generates latent environment indices with the help of the Viterbi Algorithm (Song et al.,, 2023; Elliott et al.,, 2012). To optimize the autoregressive hidden Markov model, we need to maximize its free energy lower bound, which is shown as follows:
@@ 4.3 Model Summary
Considering that the autoregressive hidden Markov model converges much faster than the sequential variational inference model, we employ a two-phase training strategy. Specifically, we first minimize $\mathcal{L}_{HMM}$ to train the autoregressive hidden Markov model. Then we minimize $\mathcal{L}_{ELBO}$ by fixing the parameters of the autoregressive hidden Markov model. Since we use the historical observations ${\mathbf{x}}_{1:t}$ to generate $\hat{{\mathbf{e}}}_{1:t}$, which can be used to estimate the transition matrix $\hat{\mathbf{A}}$, during testing phase, we can estimate $\hat{{\mathbf{e}}}_{t+1:T}$ by sampling from $\hat{\mathbf{A}}$ as shown in the environment prediction block in Figure 3. Please refer to Appendix G for model efficiency comparison.
@@ 5.1.1 Experimental Setup
Data Generation. We generate the simulated nonstationary time series data with 3 environments. Specifically, we first randomly initialize a Markov Chain with a transition matrix A. Sequentially, for each environment, we consider different Gaussian distributions and generate the nonstationary latent variables ${\mathbf{z}}_{t}^{e}$. As for the stationary latent variables, we employ an MLP with a LeakyReLU unit as the transition function. We generate Dataset A and Dataset B with different time lag dependencies. Finally, we use a randomly initialized MLP to generate the observation data.
Evaluation Metrics. We consider three different metrics to evaluate the effectiveness of our method. First, to evaluate the identifiability of the stationary and nonstationary latent variables, we consider the Mean Correlation Coefficient (MCC) on the test dataset, which is a standard metric for nonlinear ICA. A higher MCC denotes that the model can achieve better identification performance. Second, to evaluate the identifiability of the transition matrix A, we further consider the Mean Square Error (MSE) between the ground truth A and the estimated one. A lower value of MSE implies the model can identify the transition matrix better. Finally, we also consider the accuracy of ${\mathbf{e}}_{t}$ estimation since it reflects the performance of our model in detecting when the temporal distribution shift occurs. Please refer to Appendix F.1 for a detailed discussion about evaluation metrics.
Baselines. Besides the standard BetaVAE (Higgins et al.,, 2016) that does not consider any temporal and environment information, we also take some conventional nonlinear ICA methods into account like TCA (Hyvarinen and Morioka,, 2016)and i-VAE (Khemakhem et al., 2020a, ). Moreover, we consider TDRL (Yao et al.,, 2022), which considers stationary and nonstationary causal dynamics, but requires observed environment variables. Finally, we consider the HMNLICL (Hälvä and Hyvarinen,, 2020), NCTRL (Song et al.,, 2023), and CtrlNS (Song et al.,, 2024), which are designed for unobserved nonstationary.
@@ 5.1.2 Results and Discussion
Table 2: Experiment results of two synthetic datasets on baselines and proposed UDA.
@@
BetaVAE (Higgins et al.,, 2016) 64.2 63.2 63.7
TCL (Hyvarinen and Morioka,, 2016) 56.0 65.8 60.9
i-VAE (Khemakhem et al., 2020a, ) 76.9 73.0 74.9
HMNLICA (Hälvä and Hyvarinen,, 2020) 83.2 74.5 78.8
TDRL (Yao et al.,, 2022) 78.5 78.8 78.6
NCTRL (Song et al.,, 2023) 81.4 79.4 80.4
CtrlNS (Song et al.,, 2024) 87.9 85.1 86.5
UDA 97.5 92.7 95.1
@@
Experiment results of MCC are shown in Table 2, and the experiment results of environment estimation accuracy and MSE can be found in Appendix F.1. We can draw the following conclusions: 1) We can find that the accuracies of environment estimation are high in both datasets. Since Dataset B contains more complex temporal relationships, the corresponding accuracy is slightly lower. 2) We can also find that the proposed UDA model can reconstruct the latent variables under unknown temporal distribution shift with ideal MCC performance, i.e., $(>0.95)$ on average. In the meanwhile, the other compared methods, which do not use historical dependency, can hardly perform well. Moreover, the TDRL, which considers the temporal causal relationship, cannot obtain an ideal MCC performance since it requires observed environments. 3) Although baselines like HMNLICA, NCTRL and CtrlNS are devised for unobserved nonstationarity, they require strong conditions and hence can not achieve ideal identification results.
@@ 5.2.1 Experiment Setup
Datasets. We conduct experiments on eight real-world benchmark datasets that are widely used in nonstationary time series forecasting: ETT (Zhou et al.,, 2021), Exchange (Lai et al.,, 2018), ILI(CDC), electricity consuming load (ECL), weather (Wetterstation), traffic and M4 (Makridakis et al.,, 2020). More detailed descriptions of the datasets can be found in Appendix F.2.1. We employ the same data preprocessing and split ratio in TimeNet (Wu et al.,, 2022). Following the same setting of TimesNet, for each forecasting window length $H$, we let the length of the lookback window be $H$. Moreover, for each dataset, we consider different forecast lengths $H\in\{48,96,144,192\}$.
@@
Baselines. We consider the following state-of-the-art deep forecasting models for time series data. First, we consider the methods for long-term forecasting including the TCN-based methods like TimesNet (Wu et al.,, 2022) and MICN (Wang et al.,, 2022), and ModernTCN (Luo and Wang,, 2024) the MLP-based methods like DLinear (Zeng et al.,, 2023), as well as the recently proposed WITRAN (Jia et al.,, 2023), iTransformer (Liu et al., 2023a, ), and FITS (Xu et al.,, 2023). Moreover, we further consider the methods with the assumption that the temporal distribution shift occurs among instances like RevIN (Kim et al.,, 2021) and Nonstationary Transformer (Liu et al.,, 2022). Finally, we compare the nonstationary forecasting methods with the assumption that temporal distribution shift occurs uniformly, like Koopa (Liu et al., 2023b, ) and SAN (Liu et al., 2023c, ). We also consider the recent works for nonstationary time series forecasting like FOIL (Liu et al., 2024a, ), SOILD (Liu et al., 2024b, ), and FAN (Ye et al.,, 2024). We repeat each experiment over 3 random seeds and publish the average performance. Please refer to Appendix F.2.2 for more experiments.
Models UDA Koopa iTransformer Informer+FOIL PatchTST+SOILD DLinear+FAN TimesNet DLinear N-Transformer
Metric Length MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ECL 48 0.129 0.228 0.13 0.234 0.15 0.242 0.219 0.314 0.179 0.262 0.195 0.275 0.149 0.254 0.158 0.241 0.155 0.26
96 0.131 0.223 0.136 0.236 0.138 0.233 0.231 0.322 0.185 0.267 0.194 0.278 0.17 0.275 0.153 0.245 0.175 0.279
144 0.147 0.239 0.149 0.247 0.146 0.241 0.246 0.337 0.182 0.266 0.189 0.275 0.183 0.287 0.152 0.245 0.189 0.289
192 0.157 0.25 0.156 0.254 0.152 0.248 0.28 0.363 0.189 0.272 0.192 0.281 0.189 0.291 0.153 0.246 0.197 0.298
ETTh2 48 0.225 0.298 0.226 0.3 0.244 0.314 0.258 0.407 0.248 0.32 0.344 0.373 0.241 0.319 0.226 0.305 0.318 0.375
96 0.284 0.34 0.297 0.349 0.302 0.356 0.302 0.369 0.313 0.359 0.386 0.399 0.325 0.376 0.294 0.351 0.411 0.441
144 0.312 0.365 0.333 0.381 0.353 0.389 0.335 0.456 0.358 0.385 0.421 0.424 0.374 0.408 0.353 0.397 0.48 0.469
192 0.336 0.379 0.356 0.393 0.383 0.408 0.499 0.482 0.397 0.408 0.445 0.443 0.394 0.434 0.385 0.418 0.449 0.467
Exhange 48 0.042 0.141 0.042 0.143 0.045 0.148 0.063 0.146 0.0426 0.083 0.042 0.142 0.059 0.172 0.043 0.145 0.07 0.188
96 0.086 0.205 0.083 0.207 0.094 0.219 0.142 0.274 0.09 0.222 0.088 0.21 0.12 0.255 0.084 0.22 0.171 0.296
144 0.125 0.254 0.13 0.261 0.157 0.284 0.184 0.341 0.137 0.298 0.128 0.258 0.206 0.334 0.132 0.254 0.35 0.416
192 0.164 0.296 0.184 0.309 0.214 0.336 0.236 0.369 0.177 0.331 0.171 0.297 0.377 0.463 0.178 0.299 0.566 0.573
ILI 24 1.456 0.778 1.621 0.800 2.422 1.018 2.184 0.806 1.477 0.858 1.556 0.784 2.464 1.039 2.624 1.118 2.565 1.018
36 1.79 0.839 1.803 0.855 2.491 1.050 2.956 1.121 1.873 0.909 1.832 0.859 2.388 1.007 2.693 1.156 1.997 0.951
48 1.746 0.885 1.768 0.903 2.353 1.049 2.570 1.188 1.914 1.038 1.788 0.899 2.370 1.040 2.852 1.229 2.165 0.988
60 1.831 0.89 1.743 0.891 2.542 1.105 2.635 1.109 1.976 0.995 1.909 0.893 2.193 1.003 2.554 1.144 2.163 1.049
Traffic 48 0.36 0.231 0.415 0.274 0.372 0.249 0.522 0.29 0.432 0.354 0.532 0.414 0.567 0.306 0.488 0.352 0.541 0.35
96 0.323 0.218 0.401 0.275 0.337 0.233 0.53 0.293 0.401 0.325 0.476 0.375 0.611 0.337 0.485 0.336 0.529 0.349
144 0.315 0.217 0.397 0.276 0.323 0.229 0.558 0.305 0.389 0.313 0.431 0.346 0.603 0.322 0.452 0.317 0.538 0.353
192 0.315 0.22 0.403 0.284 0.322 0.233 0.589 0.328 0.399 0.308 0.434 0.348 0.604 0.321 0.438 0.309 0.507 0.342
Weather 48 0.124 0.167 0.126 0.168 0.14 0.179 0.177 0.218 0.148 0.188 0.158 0.217 0.138 0.191 0.156 0.198 0.143 0.195
96 0.151 0.301 0.154 0.205 0.168 0.214 0.225 0.259 0.187 0.226 0.199 0.265 0.18 0.231 0.186 0.229 0.199 0.246
144 0.177 0.22 0.172 0.225 0.184 0.232 0.278 0.297 0.207 0.242 0.213 0.274 0.19 0.244 0.199 0.244 0.225 0.267
192 0.193 0.233 0.193 0.241 0.203 0.252 0.354 0.348 0.234 0.265 0.238 0.298 0.212 0.265 0.217 0.261 0.296 0.315